OpenStudy (ttop0816):

PLEASE HELP! :( If |u| = 10, |v| = 8, and the angle formed between them is 60°, then u · v = ??

OpenStudy (anonymous):

You have formula to find it, it says \[cos \theta = \dfrac{u\bullet v}{|u|*|v|} \\u\bullet v= cos\theta *|u|*|v|\]

OpenStudy (ttop0816):

@Bebong is it 40 then??